A Bibliography of Publications about ANSI Standards for Programming Languages

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

31 January 2019 Version 1.27

Title word cross-reference

090 [AC88b].

1 [Nat95c, Spe79]. 1-ANSI [Spe79]. 10 [Nat95b]. 10-4 [Nat95b]. 1003.0-1995 [IEE95a]. 1003.10-1995 [IEE95b]. 1003.1c [IEE95c]. 1003.1c-1995 [IEE95c]. 1003.1I [IEE95d]. 1003.1I-1995 [IEE95d]. 1003.2-1993 [IEE93a]. 1003.2d [IEE94]. 1003.2d-1994 [IEE94]. 1003.3-1991 [IEE91]. 1003.5-1992 [IEE92]. 1003.5b [IEE96a]. 1003.5b-1996 [IEE96a]. 1003.9-1993 [IEE93b]. 102 [AC92c, AC94d]. 112 [Nat85a]. 113 [Nat85b]. 11730 [AC94a]. 11730-1994 [AC94a]. 1387.2-1995 [IEE95e].

1387.3-1996 [IEE96b]. 14.10 [AC94d]. 15 [AC88c]. 1539-1 [Ame97c]. 158 [AC88b]. 180 [Nat95c]. 180-1 [Nat95c]. 181 [Nat93]. 1815A [UA83a, UA83b, Ame95a, Uni83]. 1815A-1983 [Ame95a]. 186 [Nat00]. 186-2 [Nat00]. 1966 [Ame66, AB66, Ame78c, Ame78d]. 1968 [AC74b, AI74, AI78]. 1974 [AC74a, Ame74b, AI78, ACS85]. 1976 [AC76, AC81, Spe79, Spe81]. 1977 [AC77, AC80]. 1978 [Ame87b, AC87c]. 1980 [AC87b]. 1981 [AC87a]. 1983 [Ame83, Ame95a, UA83a, UA83b]. 1985 [Ame85d, AC90b, AC90d]. 1986 [Ame86]. 1987 [Ame87a, AC87a]. 1988 [AC88a, AA88]. 1989 [AC89a, AC89b, AN89, Ame89c, AC90a, AC90b, AC90d, ACII92]. 1989-1978 [AI78].

1989-1985 [ACS85]. 198X [AC87c, Ame87c]. 198X. [AC87d, AC86]. 1990 [AM90, AC90c, ACII92, AMM95]. 1992 [AC92a, AC92b]. 1993 [AN93, AC94b]. 1994 [AC94c, AC94a, AI94, AN94a, AN94b, AI96a]. 1995 [Ame95b, Ame95a, AIC95, AMM95, AN95a, AN95b, Nat95a]. 1996 [AI96b, AN96a, AN96b]. 1997 [Ame97c, AN97a, AN97b, AN97c, AN97d, Nat97]. 1998 [Ame98a]. 1999 [Ame99, AN99, Nat99a]. 199x [Ame90, AC92c, AC94d].

2 [Nat00]. 2003.1-1993 [IEE93c]. 2003.2-1996 [IEE96c]. 2010 [Ano11]. 2011 [Ame11].

3 [Nat99b]. 3/4/1991 [Ame91c]. 319 [Ame98a]. 319-1998 [Ame98a]. 381 [AC86]. 381-D [AC88c, AC86].

4 [Nat95b]. 46 [Nat99b]. 46-3 [Nat99b].

555-1988 [Ame97a]. 571 [AC87d]. 571-D [AC87d].

67 [Ame87b]. 67-R [Ame87b].

7185 [AC90c]. 7185-1990 [AC90c]. 754-1985 [IEE85]. 77 [Ell90]. 770.X3.97 [AC90c]. 770.X3.97-1983 [AC90c]. 770X3.97 [Ame83]. 770X3.97-1983 [Ame83]. 78-06-01 [Ame78e].

854-1987 [ANS87]. 8652 [Ame95b, Ame95a]. 8652-1995 [Ame95b, Ame95a]. 88-090 [AC88b]. 88-15 [AC88c]. 88-158 [AC88b]. 8x [MR87, Ame89a].

9075-3-1995 [Ame95c]. 9075-4-1996 [Ame96b]. 92-102 [AC92c]. 93-102 [AC94d]. 94 [BDDH95]. 9579-2-1993 [Ame93].

9593-1-1990 [Ame97b]. 96-0225 [Ame97d]. 9899 [ACII92, Ame99]. 9899-1990 [ACII92]. 9899-1999 [Ame99]. 9R1 [Spe81]. 9X [Ano93a].

Abstracts [AN97a]. Acceptance [Ame97a]. Access [Ame93, ANS92, IA78]. Account [IEE96b]. Ada [Ano93a, IEE92, IEE96a, Ame95a, Ame95b, Uni83, UA83a, UA83b]. Address [AN93]. Addresses [Pla89]. Administration [IEE95e, IEE96b]. Administrative [Nat95b]. adoption [FC83]. Algorithm [Ano11, Hop83]. Alphabet [AN98b]. Alphabetical [Nat99a]. Alphameric [ANS69a]. Amendment [IEE94, IEE95c, IEE95d, IEE96a, AC94b]. Amendments [AI00]. American [Ame81, AC76, Ame78c, Ame78d, AI83, ANS87, Ame88a, Ame88b, Ame92, AC93, AI00, Ame91c, FC83, Org91, ANS69a, Ame90, Bur83, Cad71, Hon76, Hop83, New71, Spe79, Spe81, Sch90, Uni83]. analise [Cad71]. Annotated [Sch90]. ANS [FG84]. ANS-10.2-1988 [AA88]. ANSEL [AN98b]. ANSI [Ame78c, Ame78d, AC80, AC81, ACS85, AC85, AC87b, AC87c, AC94b, AIC95, Ame96c, Ame66, AC74a, Ame74b, AC74b, AI74, AC76, AC77, AI78, Ame78c, Ame78d, AC80, AC81, AI83, Ame83, AM84, ACS85, AC85, Ame85d, Ame86, AC87b, Ame87a, Ame87b, AC87a, Ame87c, AC88a, AA88, AC89a, AC89b, AN89, Ame89b, Ame89c, AC90a, AC90b, AC90d, AM90, AC90c, Ame91b, Ame91a, ACII92, AC92a, AC92b, Ame92, Ame93, AN93, AC94b, AC94c, AC94a, AI94, AN94a, AN94b, Ame95c, Ame95b, Ame95a, AIC95, AMM95, AN95a, AN95b, Ame96b, AI96a, AI96b, AN96a, AN96b, Ame96c, Ame96a, Ame97a, Ame97c, Ame97b, AN97e, AN97a, AN97b, AN97c, AN97d, Ame98a, AN98b, AN98a, Ame98b, Ame99, AN99, Ame11, Ano11, Dyb96, IEE85, PB92, Rel87, Spe79, Spe81, Sak92, Sch90, Uni83, UA83a, UA83b].

ANSI/ANS [AA88]. ANSI/ANS-10.2-1988 [AA88]. ANSI/EIA [Ame97a]. ANSI/IEEE [AI83, Ame83, AC90c, IEE85]. ANSI/ISO [AC90c, ACII92, Ame97b]. ANSI/ISO/IEC [Ame93, AC94a, Ame95c, Ame95b, Ame95a, Ame96b, Ame97c, Ame99, Ame11]. ANSI/MDC [AM84, AM90, AMM95]. ANSI/MIL [Ame95a, Uni83, UA83a, UA83b]. ANSI/MIL-STD-1815A [Uni83]. ANSI/MIL-STD-1815A-1983 [UA83a, UA83b]. ANSI/NISO [Ame96c, AN89, AN93, AN94a, AN94b, AN95a, AN95b, AN96a, AN96b, Ame96c, AN97e, AN97a, AN97b, AN97c, AN97d, AN98b, AN98a, AN99]. ANSI/X3.215 [AC94c]. ANSI/X3.215-1994 [AC94c]. ANSI/X3.23b [AC94b]. ANSI/X3.23b-1993 [AC94b]. APG [Nat93]. API [IEE92, IEE93b, IEE95c, IEE95d, IEE96a]. Application [AN95b, IEE92, IEE95c, IEE95b, IEE96a, Ass86a, Ass86b]. Applications [IEE95d]. approved [AI74, AC76, AI78, Ame78c, Ame78d]. April [Ame78c, Ame78d]. APT [AC74a, AC77, AC80, AC87b, AIC95]. Architecture [Hor85]. Archives [AN97e]. area [ANS92]. Areas [Nat95b]. Arithmetic [ANS87, IEE85]. Arrangement [Nat99a]. Assessment [Col95]. August [AC76]. Authentication [Nat85b]. Automated [Nat93].

B [AI94]. balance [Str93]. Base [Ame97c]. Based [Ano11]. Basic [AB66, Ame84, Ame87a, AC89a, Ame98b, Hop83]. Batch [IEE94]. benefit [FC83]. Bibliographic [Ame78a, AN98b]. Binary [IEE85]. Binding [Ame96a, IEE92, IEE93b, IEE96a, Ame85a]. Bindings [Ame97b]. bit [Ins76].

C [Ame97d, Sak92, AC88b, AC88c, AC86, Ame88a, Ame88b, AC89b, Ame89c, AC90a, ACII92, Ame99, Ame11, IEE95c, IEE95d, Pla89, PB92, Pla92, Rel87, Sch90]. Call [Ame95c]. Call-Level [Ame95c]. Capability [Col95]. Carrier [ANS92]. category [Uni80]. CBEMA [Pla93]. CD [ANS92]. Challenge [Pla89]. Character [AN98b, Ame98b, AC89a]. Characters [ANS69a]. Clarification [ANS69b, ANS71]. CLI [Ame95c]. COBOL [Ame81, Ame74b, Ame74a, AC74b, AI74, AI78, ACS85, AC85, AC90b, AC90d, Ame91b, Ame91a, AC94b, FC83, FG84, Hon76, New71]. Coded [AN98b]. Codes [AN94b]. collision [ANS92]. Committees [Ano94]. Common [AC92c, AC94d, AI96a]. Compliance [Ame97a]. Computer [Ame85d, Ame86, Ame88c, Ame97b, Nat85b, AA82, AI83, Ame85b, Ame85a, AA88, IAJ83, Ins76, IA78]. Conformance [IEE91, IEE93c, IEE96c, AC93, Ame91c]. conforming [Ass86a, Ass86b]. consider [Str93]. Construction [AN98a]. contention [IA78]. Contribution [AN96b, Ame96c]. control [IA78]. corrected [AI78]. correction [AC94b]. Corrections [AN89]. Corrigenda [IEE95d]. Cost [FC83]. Cost-benefit [FC83]. Countries [Nat95b]. Creation [AN97b]. Criteria [AN99]. CSMA [ANS92]. CSMA/CD [ANS92]. Current [Hor85].

D [AC88c, AC87d, AC86]. Data [Bur83, Nat85b, Nat99b]. Database [Ame89b, Ame93, Ame95c, Ame96b, Ame92]. definicao [Cad71]. Definition [AN95b]. delicate [Str93]. Dependencies [Nat95b]. Description [Ame86, Ame87d, Ame85b]. Design [AN95a]. detection [ANS92]. Determination [Col95]. Devices [Nat97]. DIBOL [AC88a, AC92a]. Digital [Nat00]. Distributed [Bur83]. Divisions [Nat95b]. Doc [AC88b]. Document [Ame97d, Hor85].

Documents [AN97e]. Draft [Ame84, AC88b, AC87d, AC86, Ame87e, Ame81, Ame87c, Ame88a, Ame88b, Ame88c, Ame90, AC92c, AC94d, EA75, AC88c, Ame87b, Ame89a, Ame91c, FC83, Ame97d]. DSS [Nat00].

Edition [IEE91, AI78]. EIA [Ame97a]. Electronic [Org91]. Elements [AN95a]. embedding [AC87d]. Encryption [Nat99b]. engineering [AA88]. Environment [IEE94, IEE95a, IEE95b]. Environmental [Nat95a]. Errata [AI00]. Establishment [Ano11]. Evolution [Ano93a]. executive [Ins76]. Expect [Pan92]. Explained [MR87]. Extended [Ame90, AC92b, AN98b]. Extension [IEE95c, IEE95d, IEE96a]. extensions [Str93].

facilitate [AA82, AA88]. Fees [Pla93]. file [IA78]. Financial [Ano11]. FIPS [Nat85a, Nat85b, Nat93, Nat95b, Nat95c, Nat99b, Nat00]. Floating [ANS87, IEE85]. Floating-Point [ANS87, IEE85]. Form [AC94a]. Format [AN94a, AN97b, AN98a]. Formats [Hor85]. formulas [Ass86a]. Forth [AC94c]. Fortran [Ame66, Ame87c, MR87, AB66, Ame78c, Ame78b, Ame78d, Ame85a, Ame87b, AC92b, Ame96a, Ame97b, IEE93b, Ins76, IA78, Uni80, Ame87e, ANS69b, ANS71, Ame78e, AC87c, Ame89a, Ame90, Ame97c, Ano93b, Cad71, Ell90, Int87, Nat70]. full [Ame87a, AC89a, Ame98b]. Function [Ame91b, AC90b, AC90d]. Functional [Ame87d, Ame85b]. functions [Ins76].

general [AC81, AC87a]. general-purpose [AC81, AC87a]. Generalized [Ass86a, Ass86b, Ame85e, AC93, Ame91c]. Generator [Nat93]. GKS [Ame85b, Ame85a, Ame85d, Ame96a, BDDH95, HGDS86]. GKS-94 [BDDH95].

Graphical [Ame85d, Ame88c, Ame96a, HGDS86, Ame85a, Ame85b]. Graphics [Ame85d, Ame86, Ame88c, Ame97b, Ame85b, Ame85a]. Group [IEE96b]. Guide [IEE95a]. Guidelines [Nat95a, AN97a, AN98a, Nat97, Nat99a].

Hash [Nat95c]. Headers [AN96a]. Hierarchical [Ame88c, Ame97b]. High [Ano93b]. hypermedia [Gol91]. HyTime [Gol91].

Identifier [Ame96c, AN96b]. IEC [AC92b, Ame93, AC94a, Ame95c, Ame95b, Ame95a, Ame96b, Ame97c, Ame99, Ame11]. IEEE [AI83, ANS87, AI83, Ame83, AC90c]. impact [FC83]. Improvement [Col95]. Inc. [AC76, Ame78c, Ame78d]. included [Ame96a]. includes [Ame89b]. Increases [Pla93]. Independent [ANS87]. Indexes [AN99, Nat97]. Individual [Ame98b, AC89a]. Industrial [Ins76, IA78]. Industry [AN93, Ano11]. Information [ANS69a, Ame85e, Ame85d, Ame86, Ame88c, Ame89a, Ame89b, ANS92, Ame93, Ame95c, Ame95b, Ame96b, Ame97c, Ame97b, Ame97d, Ame98a, AI00, Ame11, Ano94, IEE91, IEE92, IEE93b, IEE93c, IEE93a, IEE94, IEE95c, IEE95e, IEE95d, IEE96a, IEE96b, IEE96c, AM84, ACS85, Ame85b, Ame85a, Ame85c, AC87b, Ame87a, AC87c, AC87a, AC88a, AC89a, AC89b, AC90a, AC90b, AC90d, AM90, AC90c, AC92a, Ame92, AC93, AC94b, AC94c, AI94, Ame95a, AIC95, AMM95, AI96a, AI96b, Ame91c, AN94a, AN95b, AN96a]. Initial [ANS69b]. Input [Ame98b, AC89a, Ins76]. input/output [Ins76]. Institute [Ame78c, Ame78d, Cad71]. Interactive [Ame88c, Ame97b]. Interchange [AN94a, AN94b, Gol91, Hor85]. Interface [AC94a, Ame95c, IEE92, IEE95c, IEE95e, IEE95d, IEE96a, IEE96b]. Interfaces [IEE92, IEE93b, IEE93c, IEE93a, IEE94, IEE96a, Bur83]. International [Ass86a, Ass86b, Ame97d, Col95, Hor85, Pla89]. Interpretations [AI00]. Intrinsic [Ame91b, AC90b, AC90d]. Introduction [Ell90, HGDS86, Vil94]. ISO [AI78, ACS85, Ame93, AC94a, Ame95c, Ame95b, Ame95a, Ame96b, Ame97c, Ame99, Ame11, Ass86a, Ass86b, AC90c, ACII92, AC92b, Ame97b, PB92]. ISO/IEC [AC92b]. Item [AN96b, Ame96c]. IV [Cad71].

July [AI78].

kernal [Ame85a]. Kernel [Ame85b, Ame85d, Ame96a, HGDS86]. Key [Ano11].

Language [Ass86a, Ass86b, Ame83, Ame85e, Ame88a, Ame88b, Ame89b, Ame90, Ame97b, Ame97d, AI00, Dyb96, IEE92, IEE93b, IEE95c, IEE95d, IEE96a, Pan92, Ame84, AC88b, AC88c, AC86, Ame87e, Ame81, AC74a, Ame74b, Ame74a, AC74b, AI74, AC76, AC77, AI78, Ame78c, Ame78b, Ame78d, AC80, AC81, AM84, ACS85, AC85, Ame85c, AC87b, Ame87b, AC87c, AC87a, AC88a, AC89b, Ame89a, Ame89c, AC90a, AC90b, AC90d, AC90c, AC92a, AC92b, Ame92, AC92c, AC93, AC94b, AI94, AC94d, Ame95a, AIC95, AI96a, AI96b, Ame97c, Ame91c, EA75, FC83, IAJ83, Spe79, Spe81, Str93, Uni83, UA83a, UA83b, Uni80, AI83]. Languages [Ame91b, Ame91a, AN94b, Ame95c, Ame95b, Ame98a, Ame98b, Ame99, Ame11, Sch90, AC87d, Ame87a, AC89a, AM90, ACII92, AC94c, AC94a, AMM95, Ame96b, Ame97c]. Latin [AN98b]. Lattice [Ano11]. Lattice-Based [Ano11]. layer [ANS92]. Letters [Nat99a]. Level [Ame95c, Ame97a, Spe81]. Libraries [AN97e]. Library [AN99, Pla92, Vil94]. linguagem [Cad71]. LISP [AC92c, AC94d, AI96a]. Local [Bur83, ANS92]. Lot [Ame97a].

M [AMM95]. Maintaining [Str93]. Management [AN98a]. manipulation [Ins76]. manual [Nat70, Uni83, UA83a, UA83b]. Manuscript [Org91]. Marketplace [Pla89]. Markup [Ass86a, Ass86b, Ame85e, Org91, AC93, Ame91c]. material [Ass86b]. Materials [AN99]. mathematical [Ass86a]. May [AI74, AI78]. MDC [AM84, AM90, AMM95]. Measuring [IEE91, IEE93c, IEE96c]. Metafile [Ame86]. method [ANS92]. Methods [IEE91, IEE93c, IEE96c]. Microfiche [AN96a]. Micropublishing [AN97c]. MIL [Ame95a]. MIL-STD-1815A [Uni83]. MIL-STD-1815A-1983 [UA83a, UA83b]. Minimal [Hop83]. Module [Ame91b, AC90b, AC90d]. Modules [Ame96b, Ame98b, AC89a]. Monolingual [AN98a]. multiple [ANS92].

N1043 [Ame97d]. National

[ANS69a, ANS87, Ame88a, Ame88b, Ame90, Bur83, Cad71, FC83, Hon76, Hop83, New71, Spe79, Spe81, Sch90, Uni83]. NCITS [Ame98a]. networks [ANS92]. NISO [Ame96c, AN89, AN93, AN94a, AN94b, AN95a, AN95b, AN96a, AN96b, Ame96c, AN97e, AN97a, AN97b, AN97c, AN97d, AN98b, AN98a, AN99, Nat95a, Nat97, Nat99a]. Number [AN97b, Ame97d, AN93]. Numerals [Nat99a].

Office

[Hor85, Ame85e, AC93, Ame91c, Hor85]. Open [IEE95a]. Operating [IEE93a, IEE94, IEE95c, IEE95e, IEE95d, IEE96b]. Organization [AN95a, Ano94]. Other [Nat99a]. output [Ins76]. Overview [BDDH95, Bur83].

PANCM [Ame85c]. Paper [AN97e, Ame97d, Nat95a]. Parallel [Pan92]. Part [ANS92, Ame93, Ame95c, Ame96b, Ame97c, Ame97b, IEE92, IEE93b, IEE93a, IEE94, IEE95e, IEE95d, IEE96a, IEE96b, IEE96c]. Pascal [AI83, AC90c, Ame83, IAJ83]. Password [Nat85a, Nat93]. PBASIC [Hop83]. pelo [Cad71]. Performance [Ano93b]. Permanence [AN97e]. Persistent [Ame96b]. PFORT [Ryd74]. PHIGS [Ame87d, Ame97b]. physical [ANS92]. Picture [Ame86]. PL [AC76, AC81, AC87a, AI94, EA75, Spe79, Spe81]. PL/1 [EA75, Spe79]. PL/1-ANSI [Spe79]. PL/B [AI94]. PL/I [AC76, AC81, AC87a, Spe81]. Point [ANS87, IEE85]. Polynomial [Ano11]. portability [AA82, AA88]. Portable [IEE93a, IEE94, IEE95c, IEE95e, IEE95d, IEE96b]. POSIX [IEE91, IEE92, IEE93b, IEE93c, IEE93a, IEE94, IEE95a, IEE95c, IEE95e, IEE95b, IEE95d, IEE96a, IEE96b, IEE96c]. PPM [Ame97a]. practices [AA82, AA88]. Preparation [Org91]. Preparing [Rel87]. Presentation [ANS69a]. Price [AN99]. Principal [Nat95b]. Print [AN99]. Printed [AN97d]. Procedure [Ame97a]. procedures [Ins76, IA78]. Process [Col95, Ins76]. Processing [ANS69a, Ame97b, Ame85e, ANS92]. Product [AN97c]. Profile [IEE95b]. Program [Hop83, IEE92, IEE95c, IEE95d, IEE96a]. Programmer [Ame88c, Ame97b, PB92, Spe79, Spe81]. Programming [Ame83, Ame88a, Ame88b, Ame89c, Ame90, Ame91b, Ame91a, Ame95b, Ame97c, Ame97d, Ame98a, Ame98b, Ame99, AI00, Ame11, Dyb96, Ell90, Sch90, Ame84, AC88b, AC88c, AC87d, AC86, Ame87e, Ame81, AC74a, Ame74b, Ame74a, AC74b, AI74, AC76, AC77, AI78, Ame78c, Ame78b, Ame78d, AC80, AC81, AA82, AI83, AM84, ACS85, AC85, Ame85c, AC87b, Ame87a, Ame87b, AC87c, AC87a, AC88a, AA88, AC89a, AC89b, Ame89a, AC90a, AC90b, AC90d, AM90, AC90c, ACII92, AC92a, AC92b, AC92c, AC94b, AC94c, AC94a, AI94, AC94d, Ame95a, AIC95, AMM95, AI96a, AI96b, EA75, FC83, FG84, Hon76, IAJ83, Nat70, New71, Spe79, Spe81, Uni83, UA83a, UA83b, Uni80]. programs [AA82, AA88]. Progress [ANS69b, Sak92]. Project [AC88c, AC87d, Ame87b, Col95, AC86]. Proof [AN89]. proposal [Ame88c]. Proposed [ANS69a, Ame87c, Ame88a, Ame88b, Ame90, Ame97d, Ame91c, Ame84, AC88b, AC88c, AC87d, AC86, Ame87e, Ame81, Ame87b, AC92c, AC94d, Bur83, EA75, FC83]. proposta [Cad71]. Protocol [AN95b]. PSM [Ame96b]. PUB [Nat85a, Nat85b, Nat93, Nat95b, Nat95c, Nat99b, Nat00]. Public [Ano11]. Publications [AN97e]. Publishing [AN93]. purpose [AC81, AC87a].

Quality [Ame97a].

R [Ame87b, IEE96b]. R1990 [AC90c].

R1991 [Ame91b, Ame91a]. R1993 [AIC95]. R1996 [Ame96a]. R1997 [AN97e, Ame97a, Ame97b]. R1998 [AN98b, AN98a, Ame98b]. Radix [ANS87]. Radix-Independent [ANS87]. Rationale [AC88c]. Real [IEE95d]. Real-time [IEE95d]. Realtime [IEE96a]. recommended [AA82, AA88]. Records [Nat95a]. redesignation [AC90c, ACII92, Ame95a]. Ref [AC88b]. Reference [UA83b, PB92, Spe79, Spe81, Uni83, UA83a]. References [Ame78a]. Related [Nat97]. Remote [Ame93]. Report [ANS71, AN97b]. Reports [AN95a]. Representation [AN94b]. Retrieval [AN95b, Nat97]. revised [Ame87e, Ame81, FC83]. Revision [AC74b, Ame78c, Ame78d, ACS85, AC85, AC87c, AC87a, Ame87d, AC90c, Ame95a, AMM95, AI74, AI78, AC80, AC87b, Ame87b, ACII92, Ame96c, AIC95]. REXX [AI96b, AI00].

S8 [AC87c]. SAN [AN93]. Scheme [Dyb96, Dyb96]. scientific [AA82, AA88, AN95a]. Second [ANS71]. Secure [Nat95c]. sense [ANS92]. Serial [Ame96c, AN96b]. Service [AN95b]. Services [Ano11]. Set [AN98b]. SGML [Ass86a, Ass86b, Ame85e, AC93, Ame91c]. Shell [IEE93a, IEE94, IEE96c]. Should [Pan92]. SICI [Ame96c, AN96b]. Signature [Nat00]. Smalltalk [Ame98a]. Software [Col95, IEE95e, Uni80]. Sorting [Nat99a]. Sovereignty [Nat95b]. Special [Nat95b]. Specialization [Ame93]. Specification [AN95b]. specifications [ANS92]. Specified [Ame97a]. SPICE [Col95]. Spines [AN97d]. SQL [AC87d, Ame89b, Ame92, Ame93, Ame95c, Ame96b, Ame97a]. SQL/CLI [Ame95c]. SQL/PSM [Ame96b]. Standard [Ass86a, Ass86b, AB66, AI83, Ame85e, ANS87, Ame88a, Ame88b, AC93, Ame97d, AI00, Ame91c, Bur83, Col95, Hop83, IEE85, IEE94, Nat70, Org91, Nat85a, Nat85b, Nat93, Nat95b, Nat95c, Nat99b, Nat00, New71, PB92, Sch90, Ame92, Gol91, IAJ83, IA78, Uni80, Pla89, Ame84, AC88b, AC88c, AC87d, AC86, Ame87e, Ame81, ANS69a, AC74a, Ame74b, Ame74a, AC74b, AI74, AC76, AC77, AI78, Ame78c, Ame78b, Ame78d, Ame78a, AC80, AC81, AA82, AM84, ACS85, AC85, Ame85b, Ame85a, Ame85c, AC87b, Ame87a, Ame87b, AC87c, AC87a, AC88a, AA88, AC89a, AC89b, Ame89a, AC90a, AC90b, AC90d, AM90, AC90c, Ame90, ACII92, AC92a, AC92b, AC92c, AN93, AC94b, AC94c, AC94a, AI94, AC94d, Ame95a, AIC95, AMM95, AI96a, AI96b, AN97b, EA75, Ell90, FC83, Hon76, IAJ83, Pla92, Spe79, Spe81, Uni83].

Standard [UA83a, Vil94]. Standardization [Hor85, Int87]. Standards [ANS69b, ANS71, Cad71, Gol91, Pan92, Pla93, Ano94]. statements [AC87d]. Status [Hor85]. STD [Uni83, UA83a, UA83b, ANS87]. Storage [Ame86, Nat95a]. Stored [Ame96b]. Structured [FG84, Gol91]. study [FC83]. subcategory [Uni80]. subset [AC81, AC87a]. Supercomputing [IEE95b]. Supplement [AC89a, AC90b, AC90d]. Symbols [Nat99a]. System [Ame85d, Ame88c, Ame96a, Ame97b, HGDS86, IEE92, IEE93b, IEE93c, IEE93a, IEE94, IEE95a, IEE95c, IEE95e, IEE95d, IEE96a, IEE96b, Ame85a, Ins76, IA78, Ame85b, AC94a]. Systems [Ame85d, Ame86, Ame88c, Ame89a, Ame89b, Ame97b, Ame97d, AM84, ACS85, Ame85b, Ame85a, Ame85c, Ame85e, AC87b, Ame87a, AC87c, AC87a, AC88a, AC89a, AC89b, AC90a, AC90b, AC90d, AM90, AC92a, Ame92, ANS92, AC93, AC94b, AC94c, AMM95, Ame91c].

tabular [Ass86b]. Technical

[AN95a, AN97b, IEE95d]. Technology [Ame93, Ame95c, Ame95b, Ame96b, Ame98a, Ame11, IEE91, IEE92, IEE93b, IEE93c, IEE93a, IEE94, IEE95c, IEE95e, IEE95d, IEE96a, IEE96b, IEE96c, AC90c, AC93, AC94a, Ame95a, Ame97c, Ame91c]. Template [Vil94]. Test [IEE91, IEE93c, IEE96c, Ame91c]. testing [AC93, Ame91c]. text [Ame85e, AC93]. Their [Nat95b]. Thesauri [AN98a]. Threads [IEE95c]. time [IEE95d]. TR [Nat97, Nat99a]. TR-02-1997 [Nat97]. TR-03-1999 [Nat99a]. TR01 [Nat95a]. TR01-1995 [Nat95a]. Transfer [Ame86].

unofficial [Ame78e]. update [Sak92]. Usage [Nat85a]. Use [AN98b]. User [IEE96b]. Utilities [IEE93a, IEE94, IEE96c].

Verifier [Hop83, Ryd74]. Verifying [Ame97a]. Volunteers [Pla93].

WG21 [Ame97d]. WG21/N1043 [Ame97d]. Working [Ame97d]. works [Sak92].

X11.1 [AM90, AMM95, AM84]. X11.1-1984 [AM84]. X11.1-1990 [AM90, AMM95]. X11.1-1995 [AMM95]. X3 [AC86, Ame87b]. X3.10 [AB66]. X3.10-1966 [AB66]. X3.113 [Ame87a, AC89a]. X3.113-1987 [Ame87a]. X3.113-1989 [AC89a]. X3.113a [Ame98b, AC89a]. X3.113a-1989 [Ame98b, AC89a]. X3.122 [Ame86]. X3.122-1986 [Ame86]. X3.124 [Ame85d, Ame96a]. X3.124-1985 [Ame85d, Ame96a]. X3.124.1 [Ame96a]. X3.124.1-1985 [Ame96a]. X3.135 [Ame89b, Ame92]. X3.135-1992 [Ame89b, Ame92]. X3.144.1988 [Ame88c]. X3.159 [AC86, AC89b, Ame89c, AC90a, ACII92]. X3.159-1989 [AC89b, Ame89c, AC90a, ACII92]. X3.159-198x. [AC86]. X3.165 [AC88a, AC92a]. X3.165-1988 [AC88a]. X3.165-1992 [AC92a]. X3.168 [AC87d, Ame89b]. X3.168-1989 [Ame89b]. X3.168-198X. [AC87d]. X3.198 [Ame90, AC92b]. X3.198-1992 [AC92b]. X3.198-199x [Ame90]. X3.215-1994 [AC94c]. X3.226 [AC92c, AC94d, AI96a]. X3.226-1994 [AI96a]. X3.226-199x [AC92c]. X3.23 [AC74b, AI74, AI78, ACS85, AC85, Ame91a, AC94b, Ame81, AC74b, AI74, AI78, ACS85, AC90b, AC90d, FC83]. X3.23-1968 [AC74b, AI74, AI78]. X3.23-1974 [AC74b, AI74, AI78, AC85, ACS85]. X3.23-1985 [ACS85, AC85, Ame91a, AC94b, AC90b, AC90d]. X3.238 [AI94]. X3.238-1994 [AI94]. X3.23a [Ame91b, AC90b, AC90d]. X3.23a-1989 [Ame91b, AC90b, AC90d]. X3.23b-1993 [AC94b]. X3.274 [AI96b]. X3.274-1996 [AI96b]. X3.32 [Ame74b]. X3.32-1974 [Ame74b]. X3.37 [AC80, AC87b, AIC95, AC74a, AC77, AC80, AC87b, AIC95]. X3.37-1974 [AC74a]. X3.37-1977 [AC77, AC80]. X3.37-1980 [AC80, AC87b]. X3.37-1987 [AC87b, AIC95]. X3.37-1995 [AIC95]. X3.53 [AC76, AC81, Spe79, Spe81]. X3.53-1976 [AC76, AC81, Spe79, Spe81]. X3.74 [AC81, AC87a]. X3.74-1981 [AC81, AC87a]. X3.74-1987 [AC87a]. X3.9 [Ame78c, Ame78d, AC87c, Ame66, Ame78c, Ame78d, Ame87b, AC87c, Ame87c]. X3.9-1966 [Ame66, Ame78c, Ame78d]. X3.9-1978 [Ame78c, Ame78d, Ame87b, AC87c]. X3.9-198X [AC87c, Ame87c]. X3.97 [AI83]. X3.97-1983 [AI83]. X3J11 [AC88b, AC88c]. X3J11/88 [AC88b, AC88c]. X3J11/88-090 [AC88b]. X3J11/88-15 [AC88c]. X3J11/88-158 [AC88b]. X3J13 [AC92c, AC94d]. X3J13/92 [AC92c]. X3J13/92-102 [AC92c]. X3J13/93 [AC94d]. X3J13/93-102 [AC94d]. X3J16 [Ame97d]. X3J16/96 [Ame97d]. X3J16/96-0225 [Ame97d]. X3J3 [Ame78e]. X3J3/ [Ame78e]. X9.98 [Ano11]. X9.98-2010 [Ano11].

Z39.14 [AN97a]. Z39.14-1997 [AN97a]. Z39.18 [AN95a]. Z39.18-1995 [AN95a]. Z39.19 [AN98a]. Z39.19-1993 [AN98a]. Z39.2 [AN94a]. Z39.2-1994 [AN94a]. Z39.20 [AN99]. Z39.20-1999 [AN99]. Z39.22 [AN89]. Z39.22-1989 [AN89]. Z39.23 [AN97b]. Z39.23-1997 [AN97b]. Z39.26 [AN97c]. Z39.26-1997 [AN97c]. Z39.32 [AN96a]. Z39.32-1996 [AN96a]. Z39.41 [AN97d]. Z39.41-1997 [AN97d]. Z39.43 [AN93]. Z39.43-1993 [AN93]. Z39.47 [AN98b]. Z39.47-1993 [AN98b]. Z39.48 [AN97e]. Z39.48-1992 [AN97e].

Z39.50 [AN95b]. Z39.50-1995 [AN95b]. Z39.53 [AN94b]. Z39.53-1994 [AN94b]. Z39.56 [Ame96c, AN96b]. Z39.56-1991 [Ame96c]. Z39.56-1996 [Ame96c, AN96b].

References

ANSI:1982:ANS

[AA82] American National Standards Institute and American Nuclear Society. American National Standard recommended programming practices to facilitate the portability of scientific computer programs. Technical Revision ANSI/ANS-10.2-1982, American Nuclear Society, La-Grange Park, IL, USA, March 12, 1982. 4 pp. Approved March 12, 1982 by the American National Standards Institute.

ANSI:1988:ANSb

[AA88] American National Standards Institute and American Nuclear Society. American National Standard recommended programming practices to facilitate the portability of scientific and engineering computer programs: ANSI/ANS-10.2-1988. ???? ANSI/ANS-10.2-1988, American Nuclear Society, LaGrange Park, IL, USA, April 18, 1988. 5 pp. Approved April 18, 1988.

ANSI:1966:ANS

[AB66] American National Standards Institute and Business Equipment Manufacturers Association (U.S.). American National Standard Basic FORTRAN: X3.10-1966. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1966. 32 pp.

ANSI:1974:ANSa

[AC74a] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard programming language APT: ANSI X3.37-1974. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1974. 166 pp.

ANSI:1974:ANSd

[AC74b] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard programming language COBOL: ANSI X3.23-1974: Revision of X3.23-1968. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1974. various pp.

ANSI:1976:ANS

[AC76] American National Standards Institute, Committee on Computers and Information Processing X3 and Computer and Business Equipment Manufacturers Association. American National Standard programming language PL/I: approved August 9, 1976, American National Standards Institute, Inc.: ANSI X3.53-1976. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1976. 403 pp.

ANSI:1977:ANS

[AC77] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard programming language APT: ANSI X3.37-1977. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1977. 244 pp.

ANSI:1980:ANS

[AC80] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard programming language APT: ANSI X3.37-1980, revision of ANSI X3.37-1977. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1980. 282 pp.

ANSI:1981:ANS

[AC81] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard programming language PL/I general-purpose subset: ANSI X3.74-1981: subset of ANSI X3.53-1976. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1981. 43 pp.

ANSI:1985:ANSb

[AC85] American National Standards Institute. Committee X3: Information Processing Systems and Computer and Business Equipment Manufacturers Association. American National Standard programming language COBOL: ANSI X3.23-1985 (Revision of ANSI X3.23-1974). Number 21-1 in FIPS PUB. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1985. various pp. Approved September 10, 1985, American National Standards Institute, Inc.

ANSI-X3J11:1986:DPA

[AC86] American National Standards Institute. Technical Committee X3J11 and Computer and Business Equipment Manufacturers Association. Draft proposed American National Standard programming language C: X3.159-198x. X3 project: 381-D. Computer and Business Equipment Manufacturers Association, Washington, DC, USA, October 1, 1986. various pp.

ANSI:1987:ANSe

[AC87a] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information systems: programming language — PL/I generalpurpose subset: ANSI X3.74-1987: Revision of ANSI X3.74-1981. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1987. 449 pp. Approved September 4, l987.

ANSI:1987:ANSa

[AC87b] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information systems: programming language APT: ANSI X3.37-1987, revision of ANSI X3.37-1980. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, March 20, 1987. various pp. Approved March 20, 1987.

ANSI:1987:ANSd

[AC87c] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information systems: programming language Fortran: S8 (X3.9-198X): Revision of ANSI X3.9-1978. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1987. various pp.

ANSI-X3H2:1987:DPA

[AC87d] American National Standards Institute. Technical Committee X3H2. Data Base Committee X3 Information Processing Systems and Computer and Business Equipment Manufacturers Association. Draft proposed American National Standard embedding of SQL statements into programming languages: X3.168-198X. X3 Project 571-D. Computer and Business Equipment Manufacturers Association, Washington, DC, USA, 1987. 31 pp.

ANSI:1988:ANSa

[AC88a] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information systems: programming language DIBOL: ANSI X3.165-1988. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, May 3, 1988. 86 pp. ANSI X3.165-1988: approved May 3, 1988.

ANSI-X3:1988:DPA

[AC88b] American National Standards Institute. Information Processing Systems Committee X3 and Computer and Business Equipment Manufacturers Association. Draft proposed American National Standard for information systems: programming language C: X3J11/88-158: Ref. doc. X3J11/88-090. Technical report, Computer and Business Equipment Manufacturers Association, Washington, DC, USA, 1988. viii + 218 pp. December 7, 1988.

ANSI-X3:1988:RDP

[AC88c] American National Standards Institute. Information Processing Systems Committee X3 and Computer and Business Equipment Manufacturers Association. Rationale for draft proposed American National Standard for information systems: programming language C: X3J11/88-15: Project: 381-D. Technical report, available from Global Engineering Documents, Computer and Business Equipment Manufacturers Association, 1988. vi + 119 pp. 14 November 1988.

ANSI:1989:ANSa

[AC89a] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information systems: programming languages: modules and individual character input for full BA-SIC: ANSI X3.113a-1989: Supplement to ANSI X3.113-1989. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, July 26, 1989. 14 pp. Approved July 26, 1989.

ANSI:1989:ANSb

[AC89b] American National Standards Institute and Computer and Business Equipment Manufacturers Association and Secretariat. American National Standard for information systems: programming language: C: ANSI X3.159-1989. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, December 14, 1989. 219 + vi + 119 pp.

ANSI:1990:ANSa

[AC90a] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information systems: programming language — C: ANSI X3.159-1989. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1990. 219 + vi + 119 pp. Approved December 14, 1989.

ANSI:1990:ANSb

[AC90b] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information systems: programming language: intrinsic function module for COBOL: ANSI X3.23a-1989: Supplement to ANSI X3.23-1985. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1990. ii + 16 pp. Approved September 13, 1989.

ANSI:1990:ANSe

[AC90c] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information technology: programming language PASCAL: ANSI/ISO 7185-1990: Revision and redesignation of ANSI/IEEE 770.X3.97-1983 (R1990). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, revised edition, 1990. v + 78 (printed), vi + 92 (online) pp. Available in PostScript for personal use only at http://pascal.miningco.com/msub1.htm.

ANSI:1990:ANSc

[AC90d] American National Standards Institute and Computer and Business Equipment Manufacturers Association Secretariat. American National Standard for information systems: programming language: intrinsic function module for COBOL: ANSI X3.23a-1989, supplement to ANSI X3.23-1985. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1990. ii + 76 pp.

ANSI:1992:ANSb

[AC92a] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information systems: programming language: DIBOL: ANSI X3.165-1992. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1992. x + 134 pp.

ANSI:1992:ANSc

[AC92b] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for programming language, FOR-TRAN — extended: ANSI X3.198-1992: ISO/IEC 1539: 1991 (E). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, September 21, 1992. xix + 369 pp.

ANSI:1992:DPA

[AC92c] American National Standards Institute and Computer and Business Equipment Manufacturers Association. Draft proposed American National Standard programming language Common LISP: X3.226-199x: Draft 12.24, X3J13/92-102. Computer and Business Equipment Manufacturers Association, Washington, DC, USA, July 1, 1992. ???? pp. July 1, 1992.

ANSI:1993:ANS

[AC93] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information technology — text and office systems: conformance testing for standard generalized markup language (SGML) systems. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1993. ISBN ???? v + 28 pp. LCCN ????

ANSI:1994:ANSc

[AC94a] American National Standards Institute and CODASYL Form Information Management Committee. American National Standard for information technology: programming languages: Form Interface Management System: ANSI/ISO/IEC 11730-1994. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1994. xi + 508 pp. US$18.00. URL http://

webstore.ansi.org/ansidocstore/
product.asp?sku=ANSI%2FISO%
2FIEC+11730%2D1994. Approved

December 8, 1994.

ANSI:1994:ANSa

[AC94b] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information systems: programming language: correction amendment for COBOL (ANSI X3.23-1985): ANSI/X3.23b-1993. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, January 31, 1994. various pp. Approved January 31, 1994.

ANSI:1994:ANSb

[AC94c] American National Standards Institute and Computer and Business Equipment Manufacturers Association. American National Standard for information systems: programming languages: Forth: ANSI/X3.215-1994. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, March 24, 1994. vii + 211 pp. Approved March 24, 1994.

ANSI:1994:DPA

[AC94d] American National Standards Institute. Information Processing Systems Committee X3 and Computer and Business Equipment Manufacturers Association. Draft proposed American National Standard programming language Common LISP: X3.226:199x: Draft 14.10, X3J13/93-102. X3 Secretariat/CBEMA, Washington, DC, USA, January 10, 1994. ???? pp. January 10, 1994.

ANSI:1992:ANSa

[ACII92] American National Standards Institute, Computer and Business Equipment Manufacturers Association Secretariat, International Organization for Standardization, and International Electrotechnical Commission. American National Standard for programming languages: C: ANSI/ISO 9899-1990 (revision and redesignation of ANSI X3.159-1989). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1992. viii + 219 pp. Approved August 3, 1992.

ANSI:1985:ANSa

[ACS85] American National Standards Institute, Computer and Business Equipment Manufacturers Association, and Secretariat, International Organization for Standardization. American National Standard for information systems — programming language COBOL: ANSI X3.23-1985, ISO 1989-1985: Revision of ANSI X3.23-1974. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, revised edition, 1985. various pp.

ANSI:1974:ANSe

[AI74] American National Standards Institute and International Organization for Standardization. American National Standard programming language COBOL, approved May 10, 1974: ANSI X3.23-1974: revision of X3.23-1968, volume 21-1 of Federal Information Processing Standards publication. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1974. 531 pp.

ANSI:1978:ANSa

[AI78] American National Standards Institute and International Organization for Standardization. American National Standard programming language COBOL, approved May 10, 1974: ANSI X3.23-1974: ISO 1989-1978 corrected edition, July, 1978: revision of X3.23-1968, volume 21-1 of Federal Information Processing Standards publication. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, corrected edition, 1978. 531 pp.

ANSI:1983:AIX

[AI83] American National Standards Institute and Institute of Electrical and Electronics Engineers. ANSI/ IEEE 770 X3.97-1983: An American National Standard: IEEE Standard Pascal computer programming language. IEEE, New York, NY, USA, 1983. ISBN 0-471-88944-X. 128 pp. LCCN QA76.73.P2 A43 1983.

ANSI:1994:ANSd

[AI94] American National Standards Institute and Information Technology Industry Council. American National Standard for information technology: programming language PL/B: ANSI X3.238-1994. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1994. various pp. ANSI X3.238-1994: approved December 8, 1994.

ANSI:1996:ANSa

[AI96a] American National Standards Institute and Information Technology Industry Council. American National Standard for information technology: programming language — Common LISP: ANSI X3.226-1994. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1996. various pp.

ANSI:1996:ANSb

[AI96b] American National Standards Institute and Information Technology Industry Council. American National Standard for information technology: programming language REXX: ANSI X3.274-1996. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1996. iv + 167 pp.

ANSI:2000:ANS

[AI00] American National Standards Institute and Information Technology Industry Council. American National Standard for Information Technology: Programming Language REXX — Amendments, Errata and Interpretations. 1430 Broadway, New York, NY 10018, USA, 2000. 14 pp. ANSI X3.274-1996/AM 1-2000. Approved December 13, 2000. Supplement to ANSI X3.274-1996.

ANSI:1995:ANSb

[AIC95] American National Standards Institute, Information Technology Industry Council, and Computer and Business Equipment Manufacturers Association. American National Standard for information technology: programming language APT: ANSI X3.37-1995: Revision of ANSI X3.37-1987 (R1993). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1995. various pp.

ANSI:1984:ANS

[AM84] American National Standards Institute and MUMPS Development Committee. American National Standard for information systems: programming language: MUMPS: ANSI/MDC X11.1-1984. MUMPS Development Committee, Davis, CA, USA, revised edition, 1984. ISBN 0-918118-30-1. 138 pp. LCCN ????

ANSI:1990:ANSd

[AM90] American National Standards Institute and MUMPS Development Committee. American National Standard for information systems: programming languages: MUMPS: ANSI/MDC X11.1-1990. MUMPS Users’ Group, College Park, MD, USA, November 11, 1990. 8 + 86 pp.

ANSI:1966:AF

[Ame66] American National Standards Institute. ANSI Fortran X3.9-1966. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1966. ?? pp. Approved March 7, 1966 (also known as Fortran 66). See also subsequent clarifications [ANS69b] and [ANS71].

ANSI:1974:ANSc

[Ame74a] American National Standards Institute. American National Standard programming language COBOL. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1974. 543 pp.

ANSI:1974:ANSb

[Ame74b] American National Standards Institute. American National Standard programming language: COBOL: ANSI X3.32-1974. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, revised edition, 1974. 531 pp.

ANSI:1978:ANSe

[Ame78a] American National Standards Institute. American National Standard for Bibliographic References.

American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1978. ???? pp.

ANSI:1978:ANSc

[Ame78b] American National Standards Institute. American National Standard programming language, FOR-TRAN. American National Standard; ANSI X3.9-1978 CSA standard; Z243.18-1980 American National Standards Institute. American National Standard; ANSI X3.9-1978. Canadian Standard Association. CSA standard; Z243.18-1980. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, revised edition, 1978. 438 pp. Available on the World-Wide

Web at http://observer.gsfc.
nasa.gov/iteams/doc/ansi_f77.
ps, http://observer.gsfc.nasa.
gov/iteams/doc/f77.doc, http:/
/observer.gsfc.nasa.gov/iteams/
doc/f77_cov.pdf, and http://
observer.gsfc.nasa.gov/iteams/
doc/f77_doc.pdf.

ANSI:1978:ANSb

[Ame78c] American National Standards Institute. American National Standard programming language FOR-TRAN: approved April 3, 1978, American National Standards Institute, Inc.: ANSI X3.9-1978. Revision of ANSI X3.9-1966. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, revised edition, 1978. 438 pp.

ANSI:1978:ANSd

[Ame78d] American National Standards Institute. American National Standard programming language FORTRAN: approved April 3, 1978, American National Standards Institute, Inc.: ANSI X3.9-1978. Revision of ANSI X3.9-1966. World-Wide Web document., 1978. URL http://www.

fortran.com/fortran/F77_std/

rjcnf0001.html. This is an online hypertext of the Fortran 77 Standard.

ANSI:1978:FUX

[Ame78e] American National Standards Institute. Committee on Computers X3 Information Processing. Fortran 77: unofficial, X3J3/ 90.5, (78-06-01). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1978. 371 pp.

ANSI-X3J4:1981:DPA

[Ame81] American National Standards Institute. Technical Committee X3J4. Draft proposed revised X3.23 American National Standard programming language COBOL. Computer and Business Equipment Manufacturers Association, Washington, DC, USA, 1981. various pp.

ANSI:1983:PPL

[Ame83] American National Standards Institute. The Pascal Programming Language. ANSI/IEEE 770X3.97-1983. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1983. ISBN ???? ???? pp. LCCN ????

ANSI-X3:1984:DPA

[Ame84] American National Standards Institute. Committee X3: Information Processing Systems. Draft proposed American National Standard programming language BA-SIC. Computer and Business Equipment Manufacturers Association, Washington, DC, USA, June 15, 1984. 319 pp.

ANSI:1985:ANSd

[Ame85a] American National Standards Institute. American National Standard for information systems: computer graphics — graphical kernal system (GKS) FORTRAN binding. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1985. ISBN ???? 106 pp. LCCN ????

ANSI:1985:ANSc

[Ame85b] American National Standards Institute. American National Standard for information systems: computer graphics — Graphical Kernel System (GKS) functional description. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1985. ISBN ???? 268 + 106 pp. LCCN ????

ANSI:1985:ANSe

[Ame85c] American National Standards Institute. American National Standard for information systems: programming language PANCM. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1985. ???? pp.

ANSI:1985:ISC

[Ame85d] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Information Systems—Computer Graphics— Graphical Kernel System (GKS). ANSI X3.124-1985, 1985. Includes Fortran bindings to GKS.

ANSI:1985:IPT

[Ame85e] American National Standards Institute and International Organization for Standardization. Information processing: text and office systems: Standard Generalized Markup Language (SGML). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1985. ISBN ???? viii + 158 pp. LCCN ????

ANSI:1986:ISC

[Ame86] American National Standards Institute. Information Systems—Computer Graphics—Metafile for the Storage and Transfer of Picture Description Information. ANSI X3.122-1986. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1986. ???? pp.

ANSI:1987:ANSb

[Ame87a] American National Standards Institute. American National Standard for information systems: programming languages, full BASIC: ANSI X3.113-1987. Number 68-2 in FIPS PUB. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1987. 360 pp.

ANSI:1987:ANSc

[Ame87b] American National Standards Institute. American National Standard programming language, FOR-TRAN: draft proposed: revision of ANSI X3.9-1978. X3 Project 67-R. Global Engineering Documents, Washington, DC, USA, revised edition, 1987. various pp.

ANSI:1987:DPA

[Ame87c] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed ANSI Fortran X3.9-198x, September 18, 1987. See also [MR87].

ANSI:1987:PFD

[Ame87d] American National Standards Institute. PHIGS+ Functional Description, Revision 2.0. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, July 20, 1987. ?? pp.

ANSI-X3J3:1987:DPR

[Ame87e] American National Standards Institute. Technical Committee X3J3. Draft proposed revised American National Standard programming language Fortran. X3 Secretariat/CBEMA, Washington, DC, USA, version 104 edition, 1987. various pp.

ANSI:1988:DPAa

[Ame88a] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed American National Standard Programming Language C, January 11 1988.

ANSI:1988:DPAb

[Ame88b] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed American National Standard Programming Language C, October 31 1988.

ANSI:1988:ISC

[Ame88c] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Information Systems—Computer Graphics—Programmer’s Hierarchical Interactive Graphical System. Draft proposal X3.144.1988, 1988.

ANSI:1989:ANSc

[Ame89a] American National Standards Institute. American National Standard for Information Systems programming and language: Fortran 8X draft, volume 8(4) of Fortran Forum. ACM Press, New York, NY 10036, USA, May 1989. various pp. Revision of X3.9-1978.

ANSI:1989:AXI

[Ame89b] American National Standards Institute. ANSI X3.135-1992: Information Systems — Database Language — SQL (includes ANSI X3.168-1989). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1989. ISBN ???? ???? pp. LCCN ????

US$220. URL http://www.ansi.
org/.

ANSI:1989:PLC

[Ame89c] American National Standards Institute. Programming language, C: ANSI X3.159-1989. Number 160 in FIPS Publication. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1989. 119 pp.

ANSI:1990:DPA

[Ame90] American National Standards Institute, 1430 Broadway, New York, NY 10018, USA. Draft Proposed American National Standard Programming Language Fortran Extended X3.198-199x, September 24, 1990. See also [MR87].

ANSI:1991:AXRb

[Ame91a] American National Standards Institute. ANSI X3.23-1985 (R1991): Programming Languages — COBOL. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1991. ISBN ???? ???? pp. LCCN ???? US$18.00. URL http://

webstore.ansi.org/ansidocstore/
product.asp?sku=ANSI+X3%2E23%
2D1985+%28R1991%29.

ANSI:1991:AXRa

[Ame91b] American National Standards Institute. ANSI X3.23a-1989 (R1991): Programming Languages — Intrinsic Function Module for COBOL. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1991. ISBN ???? ???? pp. LCCN ???? US$18.00. URL http://

webstore.ansi.org/ansidocstore/
product.asp?sku=ANSI+X3%2E23a%
2D1989+%28R1991%29.

ANSIX3:1991:PAN

[Ame91c] American National Standards Institute Committee X3. Proposed American National Standard: information technology: test and office systems: conformance testing for standard generalized markup language (SGML) systems: draft 3/4/1991. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1991. ISBN ???? 28 pp. LCCN ????

ANSI:1992:ANSd

[Ame92] American National Standards Institute. American national standard for information systems: database language — SQL: ANSI X3.135-1992. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1992. xiv + 580 pp. Revision and consolidation of ANSI X3.135-1989 and ANSI X3.168-1989, Approved October 3, 1989.

ANSI:1993:AII

[Ame93] American National Standards Institute. ANSI/ISO/IEC 9579-2-1993: Information Technology — Remote Database Access — Part 2: SQL Specialization. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1993. ISBN ???? ???? pp. LCCN ???? US$72. URL http://www.

ansi.org/.

ANSI:1995:ANSa

[Ame95a] American National Standards Institute. American National Standard for information technology: programming language ADA: ANSI/ISO/IEC 8652-1995: Revision and redesignation of ANSI/MIL 1815A-1983. Number 119-1 in FIPS PUB. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, revised edition, April 10, 1995. various pp.

ANSI:1995:AIIb

[Ame95b] American National Standards Institute. ANSI/ISO/IEC 8652-1995: Information Technology — Programming Languages — Ada. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1995. ISBN ???? ???? pp. LCCN ???? US$18.00. URL http://

                                        webstore.ansi.org/ansidocstore/
                                        product.asp?sku=ANSI%2FISO%
                                        2FIEC+8652%2D1995.

ANSI:1995:AIIa

[Ame95c] American National Standards Institute. ANSI/ISO/IEC 9075-3-1995: Information Technology — Database Languages — SQL — Part 3: Call-Level Interface (SQL/ CLI). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1995. ISBN ???? ???? pp. LCCN ????

                                        US$110. URL http://www.ansi.
                                        org/.

ANSI:1996:AXR

[Ame96a] American National Standards Institute. ANSI X3.124.1-1985 (R1996): Graphical Kernel System (GKS) FORTRAN Binding (included in ANSI X3.124-1985). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1996. ISBN ????

???? pp. LCCN ???? US$32. URL

http://www.ansi.org/.

ANSI:1996:AII

[Ame96b] American National Standards Institute. ANSI/ISO/IEC 9075-4-1996: Information Technology — Database languages — SQL — Part 4: Persistent Stored Modules (SQL/PSM). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1996. ISBN ???? ???? pp. LCCN

???? URL http://www.ansi.org/

.

ANSI:1996:ANZc

[Ame96c] American National Standards Institute. ANSI/NISO Z39.56-1996: Serial Item and Contribution Identifier (SICI) (revision of ANSI/ NISO Z39.56-1991). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, August 14, 1996. ???? pp.

URL https://webstore.ansi.
org/ansidocstore/product.asp?
sku=ANSI%2FNISO+Z39%2E56%2D1996.

ANSI:1997:AER

[Ame97a] American National Standards Institute. ANSI/EIA 555-1988 (R1997): Lot Acceptance Procedure for Verifying Compliance with the Specified Quality Level (SQL) in PPM. EIA (CEMA), ????, 1997. ISBN ???? ???? pp. LCCN ????

US$32. URL http://www.ansi.
org/.

ANSI:1997:AIR

[Ame97b] American National Standards Institute. ANSI/ISO 9593-1-1990 (R1997): Information Processing Systems — Computer Graphics — Programmer’s Hierarchical Interactive Graphics System (PHIGS) Language Bindings (Part 1: FOR-TRAN). American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1997. ISBN ???? ???? pp. LCCN ????

US$80. URL http://www.ansi.
org/.

ANSI:1997:AII

[Ame97c] American National Standards Institute. ANSI/ISO/IEC 1539-1:1997: Information technology — Programming languages — Fortran — Part 1: Base language. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1997. ISBN ???? ???? pp. LCCN ???? US$18.00. URL http://

webstore.ansi.org/ansidocstore/
product.asp?sku=ANSI%2FISO%
2FIEC+1539%2D1%3A1997.

ANSI:1997:WPD

[Ame97d] American National Standards Institute. Working paper for Draft Proposed International Standard for Information Systems — Programming Language C++: Document number X3J16/96-0225 WG21/N1043, December

2, 1997.  URL http://www.
maths.warwick.ac.uk/c++/pub/
;   http://www.setech.com/x3.

html. This is the first freelyavailable release of the Draft ANSI C++ Standard. This document is the second draft released for public review.

ANSI:1998:ANI

[Ame98a] American National Standards Institute. ANSI NCITS 319-1998: Information Technology — Programming Languages — Smalltalk. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1998. ISBN ???? ???? pp. LCCN ???? US$18.00. URL http://

webstore.ansi.org/ansidocstore/
product.asp?sku=ANSI+NCITS+
319%2D1998.

ANSI:1998:AXR

[Ame98b] American National Standards Institute. ANSI X3.113a-1989 (R1998): Programming Languages — Modules and Individual Character Input for Full BASIC. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1998. ISBN ???? ???? pp. LCCN ???? US$18.00. URL http://

webstore.ansi.org/ansidocstore/
product.asp?sku=ANSI+X3%2E113a%
2D1989+%28R1998%29.

ANSI:1999:AII

[Ame99] American National Standards Institute. ANSI/ISO/IEC 9899-1999: Programming Languages — C. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1999. ISBN ???? ???? pp. LCCN ???? US$18.00. URL http://

webstore.ansi.org/ansidocstore/
product.asp?sku=ANSI%2FISO%
2FIEC+9899%2D1999.

ANSI:2011:AII

[Ame11] American National Standards Institute. ANSI/ISO/IEC 9899:2011: Information Technology: Programming Languages — C. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 2011. ISBN ???? ???? pp. LCCN ???? US$232.00.

URL https://webstore.ansi.
org/RecordDetail.aspx?sku=ISO%
2fIEC+9899%3a2011.

ANSI:1995:ANSc

[AMM95] American National Standards Institute, MUMPS Development Committee, and M. Technology Association. American National Standard for information systems: programming languages: M: ANSI/MDC X11.1-1995: Revision of ANSI/MDC X11.1-1990. M Technology Association, Silver Spring, MD, USA, 1995. ISBN 0-918118-40-9. xi + 115 pp. LCCN ????

ANSI:1989:ANZ

[AN89] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.22-1989: Proof Corrections. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1989. ISBN ???? ???? pp. LCCN ????

URL ????.

ANSI:1993:ANZ

[AN93] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.43-1993: Standard Address Number (SAN) for the Publishing Industry. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1993. ISBN 1-880124-14-9. 16 pp. LCCN ???? URL http:

//www.techstreet.com/cgi-bin/
pdf/free/152624/z39-43.pdf.

ANSI:1994:ANZa

[AN94a] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.2-1994: Information Interchange Format. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1994. ISBN 1-880124-08-4. 12 pp. LCCN ???? URL http:

//www.techstreet.com/cgi-bin/
pdf/free/149295/z39-2.pdf.

ANSI:1994:ANZb

[AN94b] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.53-1994: Codes for the Representation of Languages for Information Interchange. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1994. ISBN 1-880124-10-6. 18 pp. LCCN ???? URL http:

//www.techstreet.com/cgi-bin/
pdf/free/152628/z39-53.pdf.

ANSI:1995:ANSd

[AN95a] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.18-1995: Scientific & Technical Reports: Elements, Organization, & Design. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1995. ISBN 1-880124-24-6. 30 + 16 pp. LCCN ???? URL http:

                                        //www.techstreet.com/cgi-bin/
                                        pdf/free/228864/z39-18a.pdf;
                                         http://www.techstreet.com/
                                        cgi-bin/pdf/free/228865/z39-
                                        18b.pdf.

ANSI:1995:ANZ

[AN95b] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.50-1995: Information Retrieval: Application Service Definition & Protocol Specification. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1995. ISBN 1-880124-22-X. 17 pp. LCCN ???? URL http:

                                        //www.techstreet.com/cgi-bin/
                                        pdf/free/228859/z39-50a.pdf.

ANSI:1996:ANZa

[AN96a] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.32-1996: Information on Microfiche Headers. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1996. ISBN 1-880124-25-4. 14 pp. LCCN ????

URL http://www.techstreet.
com/cgi-bin/pdf/free/152622/
z39-32.pdf.

ANSI:1996:ANZb

[AN96b] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.56-1996: Serial Item & Contribution Identifier (SICI). National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1996. ISBN 1-880124-28-9. 35 pp. LCCN ???? URL http:

//www.techstreet.com/cgi-bin/
pdf/free/152629/z39-56.pdf.

ANSI:1997:ANZa

[AN97a] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.14-1997: Guidelines for Abstracts. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1997. ISBN 1-880124-31-9. 21 pp. LCCN ???? URL http:

//www.techstreet.com/cgi-bin/
pdf/free/152600/z39-14.pdf.

ANSI:1997:ANZb

[AN97b] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.23-1997: Standard Technical Report Number Format & Creation. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1997. ISBN 1-880124-30-0. 11 pp. LCCN ???? URL http:

//www.techstreet.com/cgi-bin/
pdf/free/152603/z39-23.pdf.

ANSI:1997:ANZc

[AN97c] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.26-1997: Micropublishing Product Information. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1997. ISBN 1-880124-38-6. 11 pp. LCCN ???? URL http:

//www.techstreet.com/cgi-bin/
pdf/free/152621/z39-26.pdf.

ANSI:1997:ANZd

[AN97d] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.41-1997: Printed Information of Spines. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1997. ISBN 1-880124-32-7. 15 pp. LCCN ???? URL http:

//www.techstreet.com/cgi-bin/
pdf/free/152623/z39-41.pdf.

ANSI:1997:ANZ

[AN97e] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.48-1992 (R1997): Permanence of Paper for Publications and Documents in Libraries and Archives. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1997. ISBN 1-880124-00-9. 10 pp. LCCN ???? URL http:

//www.techstreet.com/cgi-bin/
pdf/free/136497/z39-48.pdf.

ANSI:1998:ANZb

[AN98a] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.19-1993(R1998): Guidelines for the Construction, Format, and Management of Monolingual Thesauri. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1998. ISBN 1-880124-04-1. 33 pp. LCCN ???? URL http:

//www.techstreet.com/cgi-bin/
pdf/free/228866/z39-19a.pdf.

ANSI:1998:ANZa

[AN98b] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.47-1993(R1998): Extended Latin Alphabet Coded Character Set for Bibliographic Use (ANSEL). National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1998. ISBN 1-880124-02-5. 27 pp. LCCN ???? URL http:

                                        //www.techstreet.com/cgi-bin/
                                        pdf/free/152626/z39-47.pdf.

ANSI:1999:ANZ

[AN99] American National Standards Institute and National Information Standards Organization. ANSI/ NISO Z39.20-1999: Criteria for Price Indexes for Print Library Materials. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1999. ISBN 1-880124-46-7. 19 pp. LCCN ???? URL http:

                                        //www.techstreet.com/cgi-bin/
                                        pdf/free/201273/z39-20.pdf.

Anonymous:1993:EA

[Ano93a] Anonymous. Evolution of ada 9X. ACM SIGADA Ada Letters, 13(6): 66–158, November–December 1993. CODEN AALEE5. ISSN 1094-3641 (print), 1557-9476 (electronic).

Anonymous:1993:HPF

[Ano93b] Anonymous. High Performance Fortran. Scientific Programming, 2 (1-2):1–170, Spring–Summer 1993. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Anonymous:1994:SIO

[Ano94] Anonymous. Standards information. organization of Standards Committees. ACM Fortran Forum, 13(1):4–??, March 1, 1994. ISSN 1061-7264.

Anonymous:2011:AXL

[Ano11] Anonymous. ANSI X9.98-2010:

Lattice-Based Polynomial Public Key Establishment Algorithm for the Financial Services Industry. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, April 2011.

US$100. URL http://webstore.
ansi.org/RecordDetail.aspx?
sku=ANSI+X9.98-2010.

ANSI:1969:PAN

[ANS69a] ANSI. Proposed American National Standard: Presentation of alphameric characters for information processing. Communications of the ACM, 12(12):696–698, December 1969. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

ANSI:1969:CFS

[ANS69b] ANSI Subcommittee X3J3. Clarification of Fortran standards—initial progress. Communications of the ACM, 12:289–294, 1969. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Ame66, ANS71].

ANSI:1971:CFS

[ANS71] ANSI Subcommittee X3J3. Clarification of Fortran standards— second report. Communications of the ACM, 14:628–642, 1971. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). See also [Ame66, ANS69b].

ANSI:1987:AIS

[ANS87] ANSI/IEEE. ANSI/IEEE Std 854-1987: An American National Standard: IEEE Standard for Radix-Independent Floating-Point Arithmetic. IEEE, New York, NY, USA, October 5, 1987. ISBN 0-7381-1167-8. v + 14 pp. US$44.00. URL

http://ieeexplore.ieee.org/
iel1/2502/1121/00027840.pdf;
 http://ieeexplore.ieee.org/
xpl/standardstoc.jsp?isnumber=
1121&isYear=1987.  Revised

1994. INSPEC Accession Number: 3095617.

ANSI:1992:IPS

[ANS92] ANSI. Information processing systems: local area networks — Part 3. Carrier sense multiple access with collision detection (CSMA/CD) access method and physical layer specifications. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, ANSI/ IEEE Std 802.3-1990 edition. International standard ISO/IEC 8802-3, IEEE product number: SH13482 edition, 1992. ISBN 1-55937-049-1. 220 pp. LCCN TK5105.7.I54 1990.

AAP:1986:MMF

[Ass86a] Association of American Publishers, Electronic Manuscript and Project. Markup of mathematical formulas: an SGML application conforming to International Standard ISO 8879, Standard Generalized Markup Language. Electronic manuscript series / Association of American Publishers. Association of American Publishers, Dublin, OH, USA, 1986. ISBN ???? 79 pp. LCCN Z 286 E43 M36 1989.

AAP:1986:MTM

[Ass86b] Association of American Publishers, Electronic Manuscript and Project. Markup of tabular material: an SGML application conforming to International Standard ISO 8879, Standard Generalized Markup Language. Electronic manuscript series / Association of American Publishers. Association of American Publishers, Dublin, OH, USA, 1986. ISBN 1-55653-085-4. 19 + 4 + 1 pp. LCCN Z 286 E43 M37 1989.

Brodlie:1995:GO

[BDDH95] Ken W. Brodlie, Ljiljana B.

Damnjanovic, David A. Duce, and F. Robert A. Hopgood. GKS-94: An overview. IEEE Computer Graphics and Applications, 15(6): 64–71, November 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Burr:1983:OPA

[Bur83] William E. Burr. An overview of the proposed American National Standard for Local Distributed Data Interfaces. Communications of the ACM, 26(8):554–561, August 1983. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Cadete:1971:LFI

[Cad71] Maria Odete Rodrigues Cadete. A linguagem Fortran IV; analise da definiçao proposta pelo American National Standards Institute. Technical Report ????, Instituto Gulbenkian de Ciencia, Centro de Calculo Cientifico, Oeiras, Portugal, 1971. vii + 176 pp.

Coletta:1995:SPI

[Col95] A. Coletta. The SPICE project: An international standard for software process assessment, improvement and capability determination. Lecture Notes in Computer Science, 926:49–??, 1995. CODEN LNCSD9. ISSN 0302-9743.

Dybvig:1996:SPL

[Dyb96] R. Kent Dybvig. The SCHEME Programming Language: ANSI Scheme. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1996. ISBN 0-13-454646-6. xii + 248 pp. LCCN QA76.73.S34 D93 1996. US$64.00. URL http:/

                                        /www.scheme.com/tspl2d/.

ECMA-TC:1975:DPA

[EA75] European Computer Manufacturers Association. Technical Committee, TC10-PL/1 and American National Standards Institute. Draft proposed American National Standard, programming language PL/1. Computer and Business Equipment Manufacturers Association, Washington, DC, USA, 1975. vi + 380 pp.

Ellis:1990:FPI

[Ell90] T. M. R. Ellis. Fortran 77 Programming: With an Introduction to Fortran 90 Standard. International Computer Science Series. Addison-Wesley, Reading, MA, USA, second edition, 1990. ISBN 0-201-41638-7 (paperback). xxi + 641 pp. LCCN QA76.73.F25 E43 1990. US$38.75. URL http:

//www.cbooks.com/sqlnut/SP/
search/gtsumt?source=&isbn=0-

201-41638-7. Rev. ed. of: A structured approach to FORTRAN 77 programming.

Fiorello:1983:CIS

[FC83] Marco R. Fiorello and John V.

Cugini. Cost-benefit impact study on the adoption of the draft proposed revised X3.23 American National Standard programming language COBOL. Technical Report NBSIR 83-2639, U.S. National Bureau of Standards, Gaithersburg, MD, USA, March 1983. 44 pp.

Fuori:1984:SAC

[FG84] William M. Fuori and Stephen Gaughran. Structured ANS COBOL programming. Prentice-Hall, Upper Saddle River, NJ 07458, USA, second edition, 1984. ISBN 0-13-854430-1 (paperback). x + 468 pp. LCCN ????

Goldfarb:1991:SHS

[Gol91] Charles F. Goldfarb. Standards:

HyTime: a standard for structured hypermedia interchange. Computer, 24(8):81–84, August 1991. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hopgood:1986:IGK

[HGDS86] F. Robert A. Hopgood, Julian R. Gallop, David A. Duce, and Dale C. Sutcliffe. Introduction to the Graphical Kernel System (GKS). Academic Press, New York, NY, USA, second edition, 1986. ISBN 0-12-355571-X (paperback). xii + 250 pp. LCCN T385 .I571 1986. UK£12.50. Revised for the International Standard. A. P. I. C. Studies in Data Processing No. 28.

HIS:1976:CPA

[Hon76] Honeywell Information Systems, Inc., Wellesley Hills, MA, USA. COBOL programming: American National Standard, 1976. 3 books, 51 control sheets, 2 filmstrips, 2 guides, 2 sound recordings (tape cassettes).

Hopkins:1983:APV

[Hop83] T. R. Hopkins. Algorithm 605: PBASIC: a verifier program for American National Standard Minimal BASIC. ACM Transactions on Mathematical Software, 9(4): 391–394, December 1983. CODEN ACMSCU. ISSN 0098-3500.

Horak:1985:ODA

[Hor85] Wolfgang Horak. Office Document Architecture and Office Document Interchange formats: Current status of international standardization. Computer, 18(10):50–60, October 1985. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

ISA:1978:ICS

[IA78] Instrument Society of America and American National Standards Institute. Industrial computer system FORTRAN procedures for file access and the control of file contention: standard. Instrument Society of America, Pittsburgh, PA, USA, 1978. iv + 7 pp.

IEEE-PSC:1983:ISP

[IAJ83] IEEE standard Pascal computer programming language: an American National Standard, page various, 1983. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA.

IEEE:1985:AIS

[IEE85] IEEE Task P754. ANSI/IEEE 754-1985, Standard for Binary Floating-Point Arithmetic. IEEE, New York, NY, USA, August 12, 1985. ISBN 1-55937-653-8. 20 pp. US$35.00. URL

http://ieeexplore.ieee.org/
iel1/2355/1316/00030711.pdf;
 http://standards.ieee.org/
reading/ieee/std/busarch/754-
1985.pdf;   http://standards.
ieee.org/reading/ieee/std_public/
description/busarch/754-1985_
desc.html;   http://www.iec.
ch/cgi-bin/procgi.pl/www/iecwww.
p?wwwlang=E&wwwprog=cat-det.
p&wartnum=019113. Revised 1990.

A preliminary draft was published in the January 1980 issue of IEEE Computer, together with several companion articles [?, ?, ?, ?, ?, ?, ?]. The final version was republished in [?, ?]. See also [?]. Also standardized as IEC 60559 (1989-01) Binary floating-point arithmetic for microprocessor systems.

IEEE:1991:IEI

[IEE91] IEEE. IEEE 1003.3-1991 Edition: Information Technology — Test Methods for Measuring Conformance to POSIX. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN ???? ???? pp. LCCN ???? US$54. URL http:

                                        //www.ansi.org/.

IEEE:1992:IIT

[IEE92] IEEE. IEEE 1003.5-1992: Information Technology — POSIX Ada Language Interfaces — Part 1: Binding for System Application Program Interface (API). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN ???? ???? pp. LCCN ????

                                        US$110. URL http://www.ansi.
                                        org/.

IEEE:1993:ITI

[IEE93a] IEEE. IEEE 1003.2-1993: Information Technology — Portable Operating System Interfaces (POSIX)

- Part 2: Shell and Utilities. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN ???? ???? pp. LCCN ????

                                        US$325. URL http://www.ansi.
                                        org/.

IEEE:1993:IITa

[IEE93b] IEEE. IEEE 1003.9-1993: Information Technology — POSIX FORTRAN 77 Language Interfaces — Part 1: Binding for System API. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN ???? ???? pp. LCCN ???? US$72. URL http://www.

ansi.org/.

IEEE:1993:IITb

[IEE93c] IEEE. IEEE 2003.1-1993: Information Technology — Test Methods for Measuring Conformance to POSIX — System Interfaces. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN ???? ???? pp. LCCN ????

US$120. URL http://www.ansi.
org/.

IEEE:1994:ISI

[IEE94] IEEE. IEEE 1003.2d-1994: Standard for Information Technology — Portable Operating System Interfaces (POSIX) — Part 2: Shell and Utilities — Amendment: Batch Environment. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN ???? ???? pp. LCCN

???? URL http://www.ansi.org/

.

IEEE:1995:IGP

[IEE95a] IEEE. IEEE 1003.0-1995: Guide to the POSIX Open System Environment. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN ???? ???? pp. LCCN

???? URL http://www.ansi.org/

.

IEEE:1995:IPS

[IEE95b] IEEE. IEEE 1003.10-1995: POSIX Supercomputing Application Environment Profile. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN ???? ???? pp.

                                        LCCN ???? URL http://www.
                                        ansi.org/.

IEEE:1995:IITa

[IEE95c] IEEE. IEEE 1003.1c-1995: Information Technology — Portable Operating System Interface (POSIX) - System Application Program Interface (API) Amendment 2: Threads Extension (C Language). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN ???? ???? pp. LCCN ????

                                        URL http://www.ansi.org/.

IEEE:1995:ITI

[IEE95d] IEEE. IEEE 1003.1I-1995: Information Technology — Portable Operating System Interface (POSIX)

- Part 1: System Applications Program Interface (API) — Amendment: Technical Corrigenda to Real-time Extension (C Language). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN ???? ???? pp. LCCN ????

                                        URL http://www.ansi.org/.

IEEE:1995:IITb

[IEE95e] IEEE. IEEE 1387.2-1995: Information Technology — Portable Operating System Interface (POSIX) System Administration — Part 2:

Software Administration. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN ???? ???? pp. LCCN ????

US$95. URL http://www.ansi.
org/.

IEEE:1996:IITa

[IEE96a] IEEE. IEEE 1003.5b-1996: Information Technology — POSIX Ada Language Interfaces — Part 1: Binding for System Application Program Interface (API) — Amendment 1: Realtime Extension. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN ???? ???? pp. LCCN ???? US$195. URL http://www.

ansi.org/.

IEEE:1996:IITb

[IEE96b] IEEE. IEEE 1387.3-1996: Information Technology — Portable Operating System Interface (POSIX R ) System Administration — Part 3: User and Group Account Administration. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN ???? ???? pp. LCCN ???? US$75. URL http://www.

ansi.org/.

IEEE:1996:IITc

[IEE96c] IEEE. IEEE 2003.2-1996: Information Technology — Test Methods for Measuring Conformance to POSIX — Part 2: Shell and Utilities. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996.

ISBN ???? ???? pp. LCCN ????

                                        US$125. URL http://www.ansi.
                                        org/.

ISA:1976:ICSa

[Ins76] Instrument Society of America. Industrial computer system FOR-TRAN procedures for executive functions, process input/output, and bit manipulation. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, revised: February 1976 edition, 1976. ISBN 0-87664-393-4. 11 pp.

IBM:1987:FS

[Int87] International Business Machines Corporation. Fortran standardization. IBM Corporation, San Jose, CA, USA, 1987. 43 pp.

Metcalf:1987:FE

[MR87] Michael Metcalf and John Reid. Fortran 8x Explained. Clarendon Press, Oxford, UK, 1987. ISBN 0-19-853751-4 (hardcover), 0-19-853731-X (paperback). xiv + 262 pp. LCCN QA76.73.F26 M48 1987. See also [Ame87c, Ame90].

NCCL:1970:SFP

[Nat70] National Computing Centre Limited. Standard Fortran programming manual. Computer standards series. National Computing Centre, Manchester, UK, 1970. ISBN 0-85012-021-7. [3] + 152 pp. LCCN QA76.73.F25 N37. Distributed in the U.S.A. by Science Associates/International, Inc., 23 East 26th Street, New York, N.Y. 10010. Errata sheet (3 p.) inserted.

NIST:1985:FPSa

[Nat85a] National Institute of Standards and Technology. FIPS PUB 112: Standard for Password Usage. National Institute for Standards and Technology, Gaithersburg, MD, USA, May 30, 1985.

URL http://www.itl.nist.gov/
fipspubs/fip112.htm.

NIST:1985:FPSb

[Nat85b] National Institute of Standards and Technology. FIPS PUB 113: Standard for Computer Data Authentication. National Institute for Standards and Technology, Gaithersburg, MD, USA, May 30, 1985.

URL http://www.itl.nist.gov/
fipspubs/fip113.htm.

NIST:1993:FPS

[Nat93] National Institute of Standards and Technology. FIPS PUB 181: Standard for Automated Password Generator (APG). National Institute for Standards and Technology, Gaithersburg, MD, USA, October

5, 1993. URL http://www.itl.
nist.gov/fipspubs/fip181.htm.

NISO:1995:NTE

[Nat95a] National Information Standards Organization. NISO TR01-1995: Environmental Guidelines for the Storage of Paper Records. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1995. ISBN 1-880124-21-1. 26 pp. LCCN ???? URL http:

//www.techstreet.com/cgi-bin/
pdf/free/152641/tr01.pdf.

NIST:1995:FPSa

[Nat95b] National Institute of Standards and Technology. FIPS PUB 10-4: Standard for Countries, Dependencies, Areas of Special Sovereignty, and Their Principal Administrative Divisions. National Institute for Standards and Technology, Gaithersburg, MD, USA, April 1995. URL http://www.

itl.nist.gov/fipspubs/fip10-
4.htm;   http://www.nima.mil/
gns/html/FIPS/FIP10-4.html;
http://www.nima.mil/gns/html/
FIPS/FIPS_Files.html. Super-

sedes FIPS PUB 10-3 1984 February 9. See related international standards ISO 3166-1:1997, ISO 3166-2:1998, and ISO 3166-3:1999.

NIST:1995:FPSb

[Nat95c] National Institute of Standards and Technology. FIPS PUB 180-1: Secure Hash Standard. National Institute for Standards and Technology, Gaithersburg, MD, USA, April 17, 1995. URL http://www.

itl.nist.gov/fipspubs/fip180-

1.htm. Supersedes FIPS PUB 180 1993 May 11.

NISO:1997:NTG

[Nat97] National Information Standards Organization. NISO TR-02-1997: Guidelines for Indexes and Related Information Retrieval Devices. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1997. ISBN 1-880124-36-X. 59 pp. LCCN ???? URL http:

//www.techstreet.com/cgi-bin/
pdf/free/152642/tr02.pdf.

NISO:1999:NTG

[Nat99a] National Information Standards Organization. NISO TR-03-1999: Guidelines for Alphabetical Arrangement of Letters & Sorting of Numerals & Other Symbols. National Information Standards Organization, 4733 Bethesda Avenue, Suite 300, Bethesda, MD 20814, USA (Tel: +1 301 654-2512, FAX: +1 301 654-1721), 1999. ISBN 1-880124-41-6. 28 pp. LCCN ???? URL http:

//www.techstreet.com/cgi-bin/
pdf/free/200346/tr03.pdf.

NIST:1999:FPD

[Nat99b] National Institute of Standards and Technology. FIPS PUB 46-3: Data Encryption Standard (DES). National Institute for Standards and Technology, Gaithersburg, MD, USA, October 25, 1999.

URL http://www.itl.nist.gov/
fipspubs/fip186-2.pdf. super-

sedes FIPS 46-2.

NIST:2000:FPD

[Nat00] National Institute of Standards and Technology. FIPS PUB 186-2: Digital Signature Standard (DSS). National Institute for Standards and Technology, Gaithersburg, MD, USA, January 27, 2000.

URL http://www.itl.nist.gov/
fipspubs/fip186-2.pdf.

Newell:1971:ANS

[New71] John C. Newell. American National Standard COBOL programming. Holt, Rinehart and Winston, New York, NY, USA, 1971. ISBN 0-03-086312-0. 239 pp. LCCN ????

NISO:1991:EMP

[Org91] National Information Standards Organization. Electronic Manuscript Preparation and Markup: American National Standard for Electronic Manuscript Preparation and Markup. Transaction Publishers, Rutgers University, New Brunswick, NJ 08903, USA, 1991. ISBN 0-88738-945-7. xv + 167 pp. LCCN Z286.E43 E428 1991. ANSI/NISO Z39.59-1988. Approved December 1, 1988, by American National Standards Institute. Developed by The National Information Standards Organization.

Pancake:1992:WSW

[Pan92] C. M. Pancake. What should we expect from parallel language standards? The International Journal of Supercomputer Applications, 6(1):112–117, Spring 1992. CO-DEN IJSAE9. ISSN 0890-2720.

Plauger:1992:AIS

[PB92] P. J. Plauger and Jim Brodie. ANSI and ISO Standard C: programmer’s reference. Microsoft programming series. Microsoft Press, Bellevue, WA, USA, 1992. ISBN 1-55615-359-7. vii + 278 pp. LCCN ????

Plauger:1989:SCA

[Pla89] P. J. Plauger. STANDARD C addresses challenge of the international marketplace. C Users Journal, 7(1):9–??, January 1989. ISSN 0898-9788.

Plauger:1992:SCL

[Pla92] P. J. Plauger. The Standard C Library. Prentice-Hall, Upper Saddle River, NJ 07458, USA, 1992. ISBN 0-13-838012-0. xiv + 498 pp. LCCN QA76.73.C15 P563 1991.

Plauger:1993:CIF

[Pla93] P. J. Plauger. CBEMA increases fees for standards volunteers. C Users Journal, 11(1):8–??, January 1993. ISSN 0898-9788.

Relph:1987:PAC

[Rel87] Richard Relph. Preparing for ANSI C. Dr. Dobb’s Journal of Software Tools, 12(8):16–18, 20, 22–23, August 1987. CODEN DDJOEB. ISSN 0888-3076.

Ryder:1974:PV

[Ryd74] Barbara G. Ryder. The PFORT verifier. Software—Practice and Experience, 4(4):359–377, October– December 1974. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (electronic).

Saks:1992:ACU

[Sak92] D. Saks. ANSI C++ update: works in progress. C++ Report, 4(1): 11–14, January 1992. CODEN CRPTE7. ISSN 1040-6042.

Schildt:1990:AAC

[Sch90] Herbert Schildt. The Annotated ANSI C Standard: American National Standard for Programming Languages: C. Osborne/McGraw-Hill, Berkeley, CA, USA, 1990. ISBN 0-07-881952-0. xvi + 219 pp. LCCN QA76.73.C15S356 1990. US$39.95.

SRC-UD:1979:PAN

[Spe79] Sperry Rand Corporation. Univac Division. PL/1: American National Standard programming language (PL/1-ANSI X3.53-1976): programmer reference. Sperry Rand Corp., New York, NY, USA, 1979. various pp.

SRC-UD:1981:PLA

[Spe81] Sperry Rand Corporation. Univac Division. PL/I level 9R1: American National Standard programming language PL/I, ANSI X3.53-1976: programmer reference. Sperry Univac, St. Paul, MN, USA, 1981. various pp.

Stroustrup:1993:WCL

[Str93] B. P. Stroustrup. Why consider language extensions? maintaining a delicate balance. C++ Report, 5(7): 44–51, September 1993. CODEN CRPTE7. ISSN 1040-6042.

US-DOD:1983:ANS

[UA83a] United States. Dept. of Defense and American National Standards Institute. American National Standard reference manual for the Ada programming language: ANSI/MIL-STD-1815A-1983. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, February 17, 1983. various pp.

US-DOD:1983:RMA

[UA83b] United States. Dept. of Defense and American National Standards Institute. Reference manual for the Ada programming language: ANSI/MIL-STD-1815A-1983. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. CO-DEN LNCSD9. ISBN 0-387-90887-0. ISSN 0302-9743. 330 pp. LCCN QA267.A1 L43 no.155. Approved February 17, 1983.

USNBS:1980:FCS

[Uni80] United States. National Bureau of Standards. FORTRAN: category, software standard; subcategory, programming language. FIPS pub 69, National Bureau of Standards, Washington, DC, USA, 1980. 4 pp. For sale by the National Technical Information Service.

US-DOD-AJPO:1983:ANS

[Uni83] United States. Dept. of Defense.

Ada Joint Program Office. American National Standard reference manual for the Ada programming language: ANSI/MIL-STD-1815A. American National Standards Institute, 1430 Broadway, New York, NY 10018, USA, 1983. various pp.

Vilot:1994:IST

[Vil94] M. J. Vilot. An introduction to the Standard Template Library. C++ Report, 6(8):22–29, 35, October 1994. CODEN CRPTE7. ISSN 1040-6042.